# Computational Modeling for Personalized Transcranial Electrical Stimulation: Theory, Tools, and Applications

### v1

@@ 3.2 Optimization algorithms for personalized tES
Figure 2: Schematic diagram of the stimulation optimization methods. (A) Reciprocity principle: The Reciprocity Principle posits that the path of electrical current flow from a stimulation electrode placed on the scalp to a designated area within the brain is equivalent to the path of current flow in the reverse direction, from that specific brain area back to the scalp electrode. Consequently, this principle allows for the utilization of a dipole and a simulation model to determine the optimal electrode configuration for targeted stimulation. (B) Convex optimization programming: By iteratively substituting a non-convex problem with solvable convex relaxations or linearized surrogates and refining them, one can leverage convex optimizers to efficiently approach the global optimum of the original formulation. (C) Genetic algorithm: This algorithm encodes electrode configurations into the genes of each solution within a population. Genes represent the activation status of the corresponding channel. The solutions represent a stimulation pattern.During each iteration, the solutions undergo crossover and mutation processes at random, resulting in the generation of offspring. Offspring demonstrating superior performance are retained. Ultimately, we can get the optimal solution at the last iteration. (D) Deep learning: The approach utilizes a neural network in which a constant unit replaces the input. Fully connected layers are used to generate weights within the electrode layer, which features twice as many nodes as the number of electrodes. Utilizing both the electrode layer and the lead field matrix, the distribution of the electrical field and then the loss is calculated. The goal is to optimize the electrode currents that minimize the loss by training the weights of the electrode layer. The panel is adapted from [117].
@@
A variety of methods have been proposed for optimizing electrical stimulation strategy. The exhaustive search algorithm is a brute-force method which traverses all possible solutions. Its advantage lies in its universality, suitable for tACS and tTIS, even multi-pair tTIS [118], and its ability to obtain the unique, optimal solution of the tES electrode montage and the corresponding injection current [39, 40, 119, 120]. However, given the vast search space, the exhaust search algorithm is extremely slow and computationally intensive. For example, Rampersad et al. reported that they have to evaluate all 146 million candidates in order to identify the optimal tTIS strategy involving 88 electrodes [40].
@@
The reciprocity principle, first introduced by Helmholtz in 1853, describes the interchangeability of electrical sources and measurement points in a linear, time-invariant system. It states that the effect of a source at a point in the system (e.g., the brain) can be reciprocated if the source and measurement locations are swapped. Leveraging the reciprocity principle (Fig. 2 A), tES optimization can be approached by first simulating how a source at a target brain location would influence scalp electrodes, and then inverting this relationship to determine optimal electrode placements and activation patterns for effective stimulation of the target[121, 122].
@@
In practical applications, the lead field matrix encodes this forward information. Based on the reciprocity principle, one can simplify the optimization process by selecting the channel exhibiting maximum intensity to derive an optimal stimulation pattern. Fernandez et al.[123] further advanced the understanding of the reciprocity principle by presenting a unified framework encompassing related methods. They theoretically demonstrated that Least Squares (LS), Weighted Least Squares (WLS), and reciprocity-based closed-form solutions are all special cases of the extended directional maximization problem. Moreover, LS/WLS and reciprocity-based solutions represent two extremes of the intensity-focality trade-off inherent in this optimization problem. This insight not only validates the efficacy of the reciprocity principle but also provides a means to characterize the performance bounds of other approaches without requiring extensive computational effort[22, 111].
@@
Iterative convex and linear optimization solvers are widely recognized for their efficiency and effectiveness in addressing complex problems [43, 60, 66, 113]. These approaches typically involve maximizing or minimizing carefully formulated objective functions. Variants of Least Squares (LS) optimization and constrained directional maximization are among the most fundamental and widely adopted techniques in this domain. They form the computational backbone of mainstream tES optimization software, such as SimNIBS [71, 82] and ROAST [70], and are crucial for optimizing tTIS [68, 69]. For example, a typical optimization problem in tES aims to maximize the electric field intensity at a predefined orientation while satisfying practical constraints, including a total current budget, upper bounds on field strength within avoidance zones, and per-electrode current limits. In this context, Least Squares methods minimize the second-order error between the achieved and desired electric fields, subject to current constraints [42, 60, 121]. In contrast, constrained directional maximization seeks to directly maximize the target electric field intensity, also under constraints such as current limits [60, 122] or avoidance zone regulations [113, 124].
@@
When both the objective function and the constraints are convex, powerful frameworks like Disciplined Convex Programming (Fig. 2 B) offer effective and scalable solutions [125]. However, more complex or non-convex scenarios necessitate additional algorithmic strategies. For instance, the integration of branch-and-bound algorithms enables optimizing systems with a limited number of active electrodes—a non-convex problem [66]. Branch-and-bound methods enhance computational tractability by systematically partitioning the search space and pruning suboptimal regions, ultimately converging to the global optimum. The applicability of convex optimization methods can thus be significantly extended through such hybrid approaches, allowing them to accommodate increasingly complex and realistic constraints.
@@
Machine learning has also been utilized to optimize electrode montages, such as genetic algorithms (Fig. 2 C) and neural networks (Fig. 2 D). The genetic algorithm maintains a population of candidate solutions, each solution has a set of proprieties called chromosomes. The genetic algorithm encodes the active state of each electrical channel into the chromosomes of solutions, subsequently engaging in mutation and crossover processes to recombine these chromosomes. With the evaluation of electrical field by electrode configurations derived from newly generated solutions, algorithm retains those with superior performance while discarding less effective ones. Through iterative execution, the genetic algorithm converges on the optimal solution. This approach exhibits inherent flexibility in managing non-convex formulations and demonstrates remarkable adaptability in tackling objectives and constraints of increased complexity. For example, Ruffini et al. and Lee et al. employed genetic algorithm to optimize multi-focal stimulation [39, 43], and Stoupis et al. used genetic algorithms to achieve focal stimulation over the hippocampus and thalamus using two pairs of tTIS electrodes [87].
@@
Additionally, neural networks have also been employed to address this problem. Bahn et al. applied unsupervised neural networks to optimize tES and achieve multi-target stimulation based on HD-tTIS and HD-tDCS [117]. In this unsupervised neural network, there is no traditional input. Instead, a constant value is propagated through fully connected layers to generate an electrical configuration. Thereafter, a stimulation network calculates the electric field based on this configuration, from which a loss function is derived. Through this mechanism, the neural network optimizes the electrode currents using repeated backpropagation to achieve field aligned with the target.
@@
Achieving focused stimulation is a key objective of optimization. However, there is an inherent trade-off between focality and intensity in tES. A stronger intensity within the target often results in increased activation outside the target, particularly for deep brain targets. This trade-off can be managed through additional design strategies. The Weighted Least Squares (Weighted-LS) method can enhance the focus on intensity within the target area by applying a weight matrix [60]. For constrained directional maximization, implementing an extra constraint to limit power outside the target area offers a means to balance these trade-offs [111]. The basic reciprocity-based method primarily focuses on maximizing stimulation intensity without explicit consideration of focality. Dmochowski et al. [42] addressed this limitation by incorporating L1-constraints, demonstrating that naive maximization strategies can yield suboptimal stimulation patterns. By aligning spatially decorrelated scalp potentials with optimized tES current patterns, they achieved enhanced stimulation efficacy, highlighting the importance of considering spatial focality alongside intensity.
@@
Multi-objective optimization via evolutionary algorithm (MOVEA) is another natural solution for addressing the focused stimulation. Wang et al. provide a comprehensive comparison of various modulation methods and objectives based on the Pareto front [22]. This Pareto front consists of optimal solutions that meet various requirements while respecting trade-off relationships between conflicting objectives such as intensity and focality. And MOVEA is versatile and suitable for both tACS and tTIS based on high definition and two-pair systems. Furthermore, most algorithms require the definition of a preferred direction, which may not always be available due to prior knowledge or required data. An alternative approach is to optimize the field strength. MOVEA have proven to be effective in solving it, although this introduces greater complexity due to the need to calculate the norm.
3.3 Applications for optimizing tES strategy
Personalized tES montage, in contrast to traditional montage, can achieve more effective targeted neuromodulation  [112, 126]. Therefore, inverse optimization has been extensively utilized in research. Mainstream applications respectively rely on computational modeling simulation and optimization algorithm implementation (Table LABEL:tab:application2). Computational modeling has been used to determine the optimal electrode configuration. Many studies employ multiple stimulation montages in the forward model to identify the effectiveness by analyzing the distribution of the electric field in the brain. Galletta et al. developed a forward model of a patient with left frontal stroke and compared five tDCS montages used in the clinical treatment of aphasia. Mackenbach et al. simulated and demonstrated through computational modeling that increasing the inter-electrode distance in HD-tDCS enhances the electric field strength and stimulation depth [127]. In a subsequent study, they utilized this discovery to customize a personalized montage to investigate the modulating effect of HD-tDCS on upper limb rehabilitation in ischemic stroke patients [128].
Algorithmic optimization is primarily reliant on existing implementations within modeling software. Mainstream software packages such as SimNIBS and ROAST incorporate optimization algorithm pipelines. Many studies utilize optimization results from computational models to design stimulus montages, aiming to minimize individual variability and enhance stimulus efficacy. Hsu et al. investigated the impact of tDCS on sequence learning by employing ROAST to create a forward model of each subject’s MRI. They used the targeting pipeline within ROAST to generate a personalized montage, confirming that tDCS facilitation of the motor cortex relies on a sufficiently high electric field strength [129]. Guillen et al. employed the targeting pipeline within ROAST to determine optimal electrode configurations for the MIDA model [130] while investigating the most effective tDCS montage for patients with cranial defects [131].
Furthermore, algorithmic optimization plays a critical role in guiding treatment decisions. Prior research has demonstrated that lesion location significantly influences the local electric field surrounding the stimulation target, resulting in variations in electric field strength [132]. Considering that the efficacy of tES relies on sufficient electric field strength, utilizing optimized results as a treatment paradigm can minimize variability when evaluating treatment outcomes [133]. Machado et al. utilized commercial software HD-Explore to determine electrode positions and current magnitude for HD-tDCS, while examining the effects of tDCS on athletic performance and physiological-psychological responses in athletes [134]. Grover et al. achieved longer-lasting memory enhancement by using optimization results from HD-Targets (version 3.0.1, Soterix Medical) in the treatment of memory function improvement among elderly individuals [25].
Remarkably, algorithmic optimization has garnered positive feedback regarding its therapeutic efficacy. Cruijsen et al. discovered that personalized tDCS montages enhanced intracerebral electric field strength among stroke patients [133]. Rasmussen et al. observed that personalized HD-tDCS notably enhanced delayed memory in Alzheimer’s disease (AD) patients compared to conventional electrode montages [10]. Given the superior outcomes demonstrated by personalized tES in both research and treatment, the utilization of optimization algorithms to design stimulation protocols is expected to become increasingly prevalent in future research endeavors. Maximizing the electric field in the brain will minimize the impact of inter-subject differences, resulting in more reliable data for scientific research. This will facilitate the exploration of neuromodulation mechanisms and the expansion of therapeutic applications.
Table 4: Applications for inverse optimization
Reference Heathy / Diseased Subjects Report Findings
Kolmos M, et al. (2023). Trials [135] Diseased Subjects (Ischemic Stroke) The personalized tDCS group showed a significant improvement in the Fugl-Meyer Assessment of Upper Extremity score.
Khan A, et al. (2023). Brain Stimulation [136] Healthy Subjects Targeted and optimized mc-tDCS can outperform standard bipolar stimulation and lead to better control over stimulation outcomes.
Jog MA, et al. (2023). Science Report [137] Diseased Subjects (Major Depression Disorder) Serial HD-tDCS leads to neurostructural changes at a predetermined brain target in depression.
Williamson, J. N., et al. (2023). Frontiers in human neuroscience [128] Diseased Subjects (Ischemic Stroke) HD-tDCS can improve the function of the lesioned corticospinal tract and reduce the excitability of the contralesional cortico-reticulospinal tract.
Hsu, G., et al. (2023). Brain stimulation [129] Healthy Subjects The new electrode montage can achieve stronger motor cortex polarization than alternative montages.
Guillen, A., et al. (2023). Frontiers in human neurosciencen [131] Healthy Subjects Modeling studies that take advantage of the individual anatomy of skull defects can help guide tDCS practice in patients with skull defects and skull plates.
Caulfield KA, et al. (2022). Neuromodulation [138] Healthy Subjects Using reverse-calculation modeling to produce the same high electric fields in the cortex between participants may produce future tDCS treatments more effective for working memory.
Antonenko D, et al. (2022). Alzheimers Dement (N Y) [139] Healthy Subjects Improvements in working memory were associated with individually induced electric field strengths.
Grover S, et al. (2022). Nature neuroscience [25] Diseased Subjects (Memory Impairment) The aging brain’s plasticity can be selectively and sustainably enhanced through focused neuromodulation based on memory-specific cortical circuitry’s spatiospectral parameters.
Soleimani G, et al. (2022). Brain stimulation [140] Diseased Subjects (Methamphetamine Use Disorder) Our knowledge about activated brain regions during a specific task and connectivity between active brain regions can be used to resolve ambiguity about electrode locations for network-based modulation of the human brain.
van der Cruijsen J, et al. (2022). NeuroImage: Clinical [133] Diseased Subjects (Chronic Stroke) Individualized electrode configurations increased the electrical field strength at the anatomical and functional target for stroke subjects.
Klírová M, et al. (2021). Frontiers in Systems Neuroscience [126] Healthy Subjects Individualized HD $\theta$-tACS applied via the mPFC-ACC significantly improves the post-tACS verbal component of conflict-processing.
Rasmussen I D, et al. (2021). Journal of Alzheimer’s Disease [10] Diseased Subjects (AD) HD-tDCS significantly improved delayed memory in AD.
Suen PJC, et al. (2021). Eur Arch Psychiatry Clin Neurosci [141] Diseased Subjects (Depression) EF strength simulations might be associated with further behavioral changes in depressed patients.
da Silva Machado, D. G., et al. (2021). Scientific reports [134] Healthy Subjects Neither HD-tDCS nor conventional tDCS changed exercise performance and psychophysiological responses in athletes.
Muffel T, et al. (2019). Front Aging Neurosci [142] Healthy Subjects Anodal-tDCS over S1 elicits opposing effects on proprioceptive accuracy as a function of age.
@@ 3.3 Applications for optimizing tES strategy
4 Enhancing tES Modeling through Multimodal images and Brain-State Feedback
@@
Although computational models hold great promise for personalized tES, a substantial gap remains between current modeling approaches and real-world neuromodulation outcomes. Bridging this gap requires the integration of additional information to dynamically refine models and enhance their ability to regulate brain activity effectively. Model updating primarily involves two aspects. First, anatomical segmentation and anisotropic properties in the head model can be improved using multimodal neuroimaging data (e.g., MRI, DTI), while model parameters such as tissue conductivity can be adjusted based on recorded stimulation responses (e.g., EEG, fMRI). These refinements increase the accuracy of electric field simulations and improve model adaptability to individual-specific anatomical and physiological characteristics. Second, the objective functions of the inverse optimization problem can be revised to reflect brain-state-dependent targets, guided by real-time EEG or fMRI data. Previous studies have shown that the effects of tES are highly contingent on the current functional state of the brain [1, 91]. By leveraging neural signals to infer ongoing brain states, the optimization targets can be dynamically updated, enabling more precise, adaptive, and context-aware neuromodulation.

### v2

@@ 3.2 Optimization Algorithms for Personalized tES
Figure 2: Schematic diagram of the stimulation optimization methods. (A) Reciprocity principle: The Reciprocity Principle posits that the path of electrical current flow from a stimulation electrode placed on the scalp to a designated area within the brain is equivalent to the path of current flow in the reverse direction, from that specific brain area back to the scalp electrode. Consequently, this principle allows for the utilization of a dipole and a simulation model to determine the optimal electrode configuration for targeted stimulation. (B) Convex optimization programming: By iteratively substituting a non-convex problem with solvable convex relaxations or linearized surrogates and refining them, one can leverage convex optimizers to efficiently approach the global optimum of the original formulation. (C) Genetic algorithm: This algorithm encodes electrode configurations into the genes of each solution within a population. Genes represent the activation status of the corresponding channel. The solutions represent a stimulation pattern.During each iteration, the solutions undergo crossover and mutation processes at random, resulting in the generation of offspring. Offspring demonstrating superior performance are retained. Ultimately, we can get the optimal solution at the last iteration. (D) Deep learning: The approach utilizes a neural network in which a constant unit replaces the input. Fully connected layers are used to generate weights within the electrode layer, which features twice as many nodes as the number of electrodes. Utilizing both the electrode layer and the lead field matrix, the distribution of the electric field and then the loss is calculated. The goal is to optimize the electrode currents that minimize the loss by training the weights of the electrode layer. The panel is adapted from [bahn2023computational].
@@
A variety of methods have been proposed for optimizing electrical stimulation strategy. The exhaustive search algorithm is a brute-force method which traverses all possible solutions. Its advantage lies in its universality, suitable for tACS and tTIS, even multi-pair tTIS [lee2022multipair], and its ability to obtain the unique, optimal solution of the tES electrode montage and the corresponding injection current [lee2020individually, rampersad2019prospects, alekseichuk2019electric, huang2021comparison]. However, given the vast search space, the exhaust search algorithm is extremely slow and computationally intensive. For example, Rampersad et al. reported that they have to evaluate all 146 million candidates in order to identify the optimal tTIS strategy involving 88 electrodes [rampersad2019prospects].
@@
The reciprocity principle, first introduced by Helmholtz in 1853, describes the interchangeability of electrical sources and measurement points in a linear, time-invariant system. It states that the effect of a source at a point in the system (e.g., the brain) can be reciprocated if the source and measurement locations are swapped. Leveraging the reciprocity principle (Fig. 2 A), tES optimization can be approached by first simulating how a source at a target brain location would influence scalp electrodes, and then inverting this relationship to determine optimal electrode placements and activation patterns for effective stimulation of the target[fernandez2016transcranial, cancelli2016simple].
@@
In practical applications, the lead field matrix encodes this forward information. Based on the reciprocity principle, one can simplify the optimization process by selecting the channel exhibiting maximum intensity to derive an optimal stimulation pattern. Fernandez et al.[fernandez2020unification] further advanced the understanding of the reciprocity principle by presenting a unified framework encompassing related methods. They theoretically demonstrated that Least Squares (LS), Weighted Least Squares (WLS), and reciprocity-based closed-form solutions are all special cases of the extended directional maximization problem. Moreover, LS/WLS and reciprocity-based solutions represent two extremes of the intensity-focality trade-off inherent in this optimization problem. This insight not only validates the efficacy of the reciprocity principle but also provides a means to characterize the performance bounds of other approaches without requiring extensive computational effort[wang2022multi, huang2018optimized].
@@
Iterative convex and linear optimization solvers are widely recognized for their efficiency and effectiveness in addressing complex problems [dmochowski2011optimized, saturnino2019accessibility, guler2016optimization, ruffini2014optimization]. These approaches typically involve maximizing or minimizing carefully formulated objective functions. Variants of Least Squares (LS) optimization and constrained directional maximization are among the most fundamental and widely adopted techniques in this domain. They form the computational backbone of mainstream tES optimization software, such as SimNIBS [saturnino2019simnibs, saturnino2021optimizing] and ROAST [huang2019realistic], and are crucial for optimizing tTIS [huang2020optimization, huang2019can]. For example, a typical optimization problem in tES aims to maximize the electric field intensity at a predefined orientation while satisfying practical constraints, including a total current budget, upper bounds on field strength within avoidance zones, and per-electrode current limits. In this context, Least Squares methods minimize the second-order error between the achieved and desired electric fields, subject to current constraints [dmochowski2011optimized, fernandez2016transcranial, dmochowski2017optimal]. In contrast, constrained directional maximization seeks to directly maximize the target electric field intensity, also under constraints such as current limits [dmochowski2011optimized, cancelli2016simple] or avoidance zone regulations [guler2016optimization, guler2016optimizing].
@@
When both the objective function and the constraints are convex, powerful frameworks like Disciplined Convex Programming (Fig. 2 B) offer effective and scalable solutions [grant2014cvx]. However, more complex or non-convex scenarios necessitate additional algorithmic strategies. For instance, the integration of branch-and-bound algorithms enables optimizing systems with a limited number of active electrodes—a non-convex problem [saturnino2019accessibility]. Branch-and-bound methods enhance computational tractability by systematically partitioning the search space and pruning suboptimal regions, ultimately converging to the global optimum. The applicability of convex optimization methods can thus be significantly extended through such hybrid approaches, allowing them to accommodate increasingly complex and realistic constraints.
@@
Machine learning has also been utilized to optimize electrode montages, such as genetic algorithms (Fig. 2 C) and neural networks (Fig. 2 D). The genetic algorithm maintains a population of candidate solutions, each solution has a set of proprieties called chromosomes. The genetic algorithm encodes the active state of each electrical channel into the chromosomes of solutions, subsequently engaging in mutation and crossover processes to recombine these chromosomes. With the evaluation of electric field by electrode configurations derived from newly generated solutions, algorithm retains those with superior performance while discarding less effective ones. Through iterative execution, the genetic algorithm converges on the optimal solution. This approach exhibits inherent flexibility in managing non-convex formulations and demonstrates remarkable adaptability in tackling objectives and constraints of increased complexity. For example, Ruffini et al. and Lee et al. employed genetic algorithm to optimize multi-focal stimulation [ruffini2014optimization, lee2020individually], and Stoupis et al. used genetic algorithms to achieve focal stimulation over the hippocampus and thalamus using two pairs of tTIS electrodes [stoupis2022non].
@@
Additionally, neural networks have also been employed to address this problem. Bahn et al. applied unsupervised neural networks to optimize tES and achieve multi-target stimulation based on HD-tTIS and HD-tDCS [bahn2023computational]. In this unsupervised neural network, there is no traditional input. Instead, a constant value is propagated through fully connected layers to generate an electrical configuration. Thereafter, a stimulation network calculates the electric field based on this configuration, from which a loss function is derived. Through this mechanism, the neural network optimizes the electrode currents using repeated backpropagation to achieve field aligned with the target.
@@
Achieving focused stimulation is a key objective of optimization. However, there is an inherent trade-off between focality and intensity in tES. A stronger intensity within the target often results in increased activation outside the target, particularly for deep brain targets. This trade-off can be managed through additional design strategies. The Weighted Least Squares (Weighted-LS) method can enhance the focus on intensity within the target area by applying a weight matrix [dmochowski2011optimized]. For constrained directional maximization, implementing an extra constraint to limit power outside the target area offers a means to balance these trade-offs [huang2018optimized]. The basic reciprocity-based method primarily focuses on maximizing stimulation intensity without explicit consideration of focality. Dmochowski et al. [dmochowski2017optimal] addressed this limitation by incorporating L1-constraints, demonstrating that naive maximization strategies can yield suboptimal stimulation patterns. By aligning spatially decorrelated scalp potentials with optimized tES current patterns, they achieved enhanced stimulation efficacy, highlighting the importance of considering spatial focality alongside intensity.
@@
Multi-objective optimization via evolutionary algorithm (MOVEA) is another natural solution for addressing the focused stimulation. Wang et al. provide a comprehensive comparison of various modulation methods and objectives based on the Pareto front [wang2022multi]. This Pareto front consists of optimal solutions that meet various requirements while respecting trade-off relationships between conflicting objectives such as intensity and focality. And MOVEA is versatile and suitable for both tACS and tTIS based on high definition and two-pair systems. Furthermore, most algorithms require the definition of a preferred direction, which may not always be available due to prior knowledge or required data. An alternative approach is to optimize the field strength. MOVEA has proven to be effective in solving it, although this introduces greater complexity due to the need to calculate the norm.
@@
4 Enhancing tES Modeling through Multimodal Images, Neural Mass Model and Brain-State Feedback
@@
Although computational models hold great promise for personalized tES, a substantial gap remains between current modeling approaches and real-world neuromodulation outcomes. Bridging this gap requires the integration of additional information to dynamically refine models and enhance their ability to regulate brain activity effectively. Model updating primarily involves three aspects. First, anatomical segmentation and anisotropic properties in the head model can be improved using multimodal neuroimaging data (e.g., MRI, DTI). In contrast, model parameters such as tissue conductivity can be adjusted based on recorded stimulation responses (e.g., EEG, fMRI). These refinements increase the accuracy of electric field simulations and improve model adaptability to individual-specific anatomical and physiological characteristics. Second, macroscopic models can be integrated with mathematical formulations to enable multi-scale coupling of neural dynamics. By incorporating individual structural connectivity information, personalized neural dynamic models can be linked to anatomical models, capturing the real-time brain responses to external stimulation and inter-individual variability in outcomes. Third, the objective functions of the inverse optimization problem can be revised to reflect brain-state-dependent targets, guided by real-time EEG or fMRI data. Previous studies have shown that the effects of tES are highly contingent on the current functional state of the brain [Krause2023TranscranialES, luo2025frequency]. By leveraging neural signals to infer ongoing brain states, the optimization targets can be dynamically updated, enabling more precise, adaptive, and context-aware neuromodulation.
In this section, we outline three directions toward enhancing the fidelity and functional relevance of computational modeling for personalized tES. Section 4.1 discussed improvements in anatomical modeling. Section 4.2 focused on linking macroscopic anatomical models with neural dynamic formulations. Section 4.3 introduced adaptive optimization frameworks guided by real-time neural signals, where brain-state-dependent objectives dynamically adjust stimulation parameters for more precise and context-aware neuromodulation. Collectively, these advances outline a pathway toward closed-loop, data-driven, and truly personalized tES modeling and control.